# Towards Stable Machine Learning Model Retraining via Slowly Varying Sequences

### v4

@@ 3 Numerical Experiments
3.1 Case Study
Dataset.
@@ 3.1 Case Study
We collect daily patient data from a large hospital system in Connecticut. We extract patient features including demographics, patient status, clinical measurements, lab results, diagnoses, orders, procedures, and notes. Patient records are retrospectively identified from January 1st, 2018 to May 1st, 2022, overseeing 168,815 patients, and 865,954 patient-day records. We focus on the binary classification task of predicting patient mortality risk, defined as the likelihood of death or hospice care by hospital discharge. We evaluate model performance using the Area Under the Receiver Operating Characteristic Curve (AUC), which assesses the model’s ability to differentiate between outcomes by comparing the true positive rate against the false positive rate at various thresholds.
Data preprocessing.
@@
In the existing production pipeline, models are trained on a 6-month basis, although the retraining window is scheduled to be reduced to 3-4 months in 2024. There are no obvious data shifts for this case study. To simulate the process of receiving data updates, we divide the 52 months into multiple sub-intervals. We follow traditional ML training procedures by dividing the dataset into training, validation, and testing sets. Specifically, all patient records admitted from September 2021 to December 2022 (16 months) are allocated as the out-of-sample testing set.
Research questions.
@@
To assess the performance of the slowly varying methodology (SVML), we evaluate both the intra-sequence and the inter-sequence stability of the generated sequences. The former serves to illuminate the predictive power-stability trade-off that our methodology hopes to optimize, whereas the latter analyzes the stability of the adaptive model retraining strategy. Furthermore, we perform the following robustness checks: first, we investigate the sensitivity of SVML to the update frequency (or interval length); second, we examine the robustness of SVML with respect to different distance measures and the consistency of the resulting analytical insights.
Experimental methodology.
@@
We obtain experimental results as follows. Having selected an update frequency, the number of batches $B$ is fixed. For each batch $b\in B$, we define the available training data $D_{b}$ as the first $b$ intervals and reserve the $(b+1)$-th interval as a validation set. We generate a set of pre-computed candidate models, $\mathcal{F}_{b}$, for each batch, bootstrapping to ensure diversity in the candidate models. Then, for accuracy tolerances [0.1, 0.05, 0.02, 0.01], we generate a slowly varying sequence according to the methodology outlined in Section 2.3. To benchmark the performance of these sequences, we also generate a “greedy sequence” constructed of greedily selected, best-performing models (evaluated on the validation set). We denote by $f_{b}$ the batch $b$ model chosen by the slowly varying sequence and denote by $\bar{f}_{b}$ the batch $b$ model chosen by the greedy sequence. All AUC results are reported on the 16-month testing set put aside at the start of model training.
@@ 3.2 Results
Intra-sequence stability.
We start with an in-depth analysis of the distance-accuracy trade-off implicit in the slowly varying methodology. We select a 3-month update frequency, and conduct this experiment for three classes of models: logistic regression, CART decision trees, and XGBoost. For CART, we construct two sets of slowly varying sequences; one using a structural distance and the other using a feature importance-based distance. For each model type and each sequence, we graph the AUC and intra-sequence pairwise distance of the models comprising the slowly varying and greedy sequences in Fig. 2. Differences in test AUCs of the SVML and greedy sequences grow smaller as we reduce accuracy tolerance. Indeed, for the $\alpha=0.01$ SVML sequences, we observe only a 1.2% reduction in AUC for logistic regression and negligible reductions for the other model types. However, we note an obvious advantage in stability measured by pairwise distance, ranging from 5.5% (logistic regression) to 20% (XGBoost).
Figure 2: Test AUC (top) and distance between adjacent models (bottom) for various model/distance types
The intra-sequence stability results demonstrate the efficacy of SVML in minimizing the pairwise distance of adjacent models while sacrificing small model accuracy. By setting an accuracy tolerance of $0.01$, for example, we guarantee that the validation AUC of chosen slowly varying models is within $1\%$ of the maximum: a bound that generalizes well experimentally to the test AUC. As expected, the AUC of the models and the pairwise distances for both SVML and greedy sequences tend to, respectively, increase and decrease over time as they are given more data to be trained on.
Inter-sequence stability.
@@
Table 1: Inter-sequence stability loss for greedy and SVML sequences of various lengths
@@
Next, we empirically investigate the sensitivity of SVML to different update frequencies by splitting the available 52 months of data into intervals of different lengths (1, 3, 6, and 9 months). As discussed, the currently deployed 6-month frequency will change, so our experiment also serves to understand the properties of different frequencies. We choose XGBoost as the canonical model type for the experiment since this is the model in use in production, and train 25 candidate models per batch. The results, depicted in Fig. 3, demonstrate that trends and insights are consistent across the different frequencies tested. A 3-month re-training frequency exhibits more stable behavior compared to 1-month, while retaining high AUC ranging from 0.85-0.9 with improvements in distance from [0.3-0.37] to [0.275-0.21] compared to 6- or 9-month intervals.
Figure 3: AUC (top) and distance (bottom) between adjacent XGBoost models for various update frequencies.
@@
As a final robustness check, we investigate how well the analytical insights of an SVML sequence optimized for one distance measure (XGBoost with Gain importance) generalize to a different measure (SHAP) in terms of their stability across retraining iterations. Figure 4 presents the empirical distribution of the absolute deviation in features’ SHAP values between adjacent models for both greedy and SVML sequences. Greedy achieves a mean deviation of 0.086 with a standard deviation of 0.101 and a maximum deviation of 0.625; in comparison, SVML exhibits an improved mean deviation of 0.081 with a standard deviation of 0.094 and a maximum deviation of 0.606. These results indicate that SVML’s improved stability across retraining iterations effectively translates into consistent SHAP values, affirming the method’s reliability.
Figure 4: Empirical distribution of absolute deviation in SHAP between adjacent models.

### v5

@@ 5 Numerical Experiments
5.1 Case Studies
We validate our methodology through three case studies across different domains: in healthcare, we deploy the proposed methodology in production in collaboration with a large US hospital; in vision, we analyze model stability in a facial recognition task using the Yearbook dataset; and in language, we perform dynamic news categorization using the HuffPost dataset Yao et al. (2022).
@@ 5.1 Case Studies
Healthcare. We collect daily patient data from a large hospital system in Connecticut. We extract patient features including demographics, patient status, clinical measurements, lab results, diagnoses, orders, procedures, and notes. Patient records are retrospectively identified from January 1st, 2018 to May 1st, 2022, overseeing 168,815 patients, and 865,954 patient-day records. We focus on the binary classification task of predicting patient mortality risk, defined as the likelihood of death or hospice care by hospital discharge. We evaluate model performance using the Area Under the Receiver Operating Characteristic Curve (AUC), which assesses the model’s ability to differentiate between outcomes by comparing the true positive rate against the false positive rate at various thresholds.
@@
Data preprocessing. In the existing production pipeline, models are trained on a 6-month basis, although the retraining window is scheduled to be reduced to 3-4 months. There are no obvious data shifts for this case study. To simulate the process of receiving data updates, we divide the 52 months into multiple sub-intervals. We follow traditional ML training procedures by dividing the dataset into training, validation, and testing sets. Specifically, all patient records admitted from September 2021 to December 2022 (16 months) are allocated as the out-of-sample testing set.
Vision. Consisting of 37,921 American high school yearbook photos, the Yearbook dataset encompasses the evolving social norms, fashion styles and population demographics over the period of 1930-2013. Each photo is a $32\times 32\times 1$ gray-scale image associated with a binary outcome of the individual’s gender.
Data preprocessing. Similar to the approach for the healthcare dataset, we divide the full time sequence into training, validation, and testing given their time sequence. There are 84 years in total, and we consider every 7 years as a single time period. All photos from and including 2010 are allocated as the out-of-sample testing set. Instead of feeding the raw image as our input, we first process embeddings using open-source, pre-trained vision models that are finetuned on similar datasets. Specifically, we process the existing image to be compatibly shaped tensored to be passed to FaceNet, an existing Inception Residual Masking Network pretrained on VGGFace2 to classify facial identities. Each image is reduced down to an embedding of the shape 512, which are then used to fit the ML models of which we study stability tradeoffs on.
Language. The Huffpost dataset contains around 200,000 samples and aims to identify tags of news articles from their corresponding headlines, which are classified into 11 different categories (Black Voices, Business, Comedy, Crime, Entertainment, Impact, Queer Voices, Science, Sports, Tech, Travel).
Data preprocessing. The dataset ranges from 2012 to 2018, where each year is considered as a single individual time period in the experiment. All news headlines of 2018 are allocated as out-of-sample testing set. Similar to the vision dataset, instead of directly using a large language model to study their stability, we first process each sample using a pretrained model to obtain their respective embeddings. In particular, we use the BAAI FlagEmbedding model to map the original text into lower-dimensional vectors of size 2014.
5.2 Research Questions and Methodology
@@
Research questions. Our main experiment serves to illuminate the predictive power-stability trade-off that the slowly varying methodology (SVML) hopes to optimize. To assess the performance of our methodology with respect to this trade-off, we evaluate the predictive power and the intra-sequence stability of the generated sequences, and compare them against those of a greedily trained sequence. Furthermore, as secondary research questions, we perform the following robustness checks: (i) first, we analyze the stability of the adaptive model retraining strategy by looking at the inter-sequence stability of the generated sequences; (ii) second, we investigate the sensitivity of SVML to the update frequency (or interval length); (iii) third, we examine the robustness of SVML with respect to different distance measures and the consistency of the resulting analytical insights.
@@
Experimental methodology. We obtain experimental results as follows. Having selected an update frequency, the number of batches $B$ is fixed. For each batch $b\in B$, we define the available training data $D_{b}$ as the first $b$ intervals and reserve the $(b+1)$-th interval as a validation set. We generate a set of pre-computed candidate models, $\mathcal{F}_{b}$, for each batch, bootstrapping to ensure diversity in the candidate models. Then, for accuracy tolerances [0.1, 0.05, 0.02, 0.01], we generate a slowly varying sequence according to the methodology outlined in Section 3.1. To benchmark the performance of these sequences, we also generate a “greedy sequence” constructed of greedily selected, best-performing models (evaluated on the validation set). We denote by $f_{b}$ the batch $b$ model chosen by the slowly varying sequence and denote by $\bar{f}_{b}$ the batch $b$ model chosen by the greedy sequence. All AUC results are reported on the 16-month testing set put aside at the start of model training.
@@ 5.3 Results
Trading-off Predictive Power and Stability.
We start with an in-depth analysis of the distance-accuracy trade-off implicit in the slowly varying methodology. we compare the performance of our slowly varying machine learning (SVML) methodology against a greedy retraining approach across the three datasets (Yearbook, Healthcare, Huffpost) and four model families (logistic regression, decision trees, boosted trees, and neural networks). The results are summarized in Tables 1 and 2. We include a detailed presentation of the results in Appendix D.
Table 1 presents the percent improvement in stability loss $\mathcal{L}_{S}$ as a function of the accuracy tolerance $\alpha$. It quantifies the extent to which SVML reduces model instability across retraining iterations, demonstrating that greater stability is achieved as the accuracy threshold increases. In some cases, the same models were selected for different thresholds, which is expected when the thresholds were not large enough to shift the optimal solution. Table 2, on the other hand, quantifies the percent deterioration in realized out-of-sample accuracy due to applying SVML instead of a greedy retraining approach. We make the following observations:
  • On average, across all datasets and model types, a $2\%$ reduction in (realized) predictive power results in a $30\%$ improvement in model stability.
  • As the accuracy tolerance increases, stability gains become more pronounced, confirming that allowing small sacrifices in predictive power leads to smoother model updates.
  • The degree of stability improvement varies by model type, with interpretable models (logistic regression, CART) and neural networks benefiting more from stability constraints than ensemble models like XGBoost (which are known to be more stable).
  • The actual, realized deterioration in out-of-sample predictive power is always significantly smaller than the pre-selected accuracy tolerance, suggesting that stabilization often leads to improved generalization. (E.g., negative deterioration is due to the fact that the greedy methodology selects the best performing model using a validation set, so, when evaluated on the test set, its performance could be worse that SVML.)
To summarize, (i) across all models and datasets, as we allow a higher accuracy tolerance, the improvement in stability consistently increases. This is expected, as relaxing the predictive power constraint enables SVML to prioritize selecting smoother model transitions. (ii) While all models benefit from the SVML approach, the stability gains are most pronounced for logistic regression, decision trees, and neural networks. These models exhibit inherently higher sensitivity to data perturbations, and thus benefit more from explicitly enforcing stability. (iii) The average results confirm that even for small sacrifices in predictive power (1-2%), we can achieve meaningful reductions in model instability. This is important in real-world deployment settings, where sudden changes in model outputs can lead to user distrust or operational inefficiencies.
Dataset Yearbook Healthcare Huffpost Avg.
$\alpha$ 0.01 0.02 0.05 0.1 0.01 0.02 0.05 0.1 0.01 0.02 0.05 0.1 0.01 0.02 0.05 0.1
LogReg 23.76 23.76 23.76 23.76 46.67 60.86 69.84 69.84 8.74 8.74 8.74 8.74 26.39 31.12 34.11 34.11
CART 15.50 26.16 26.91 26.91 23.86 25.02 29.67 29.67 22.36 26.65 30.8 30.8 20.57 25.94 29.13 29.13
XGBoost 11.09 11.09 11.09 11.09 23.43 25.03 25.03 25.03 7.17 7.17 7.17 7.17 13.90 14.43 14.43 14.43
MLP 18.29 23.21 23.21 23.21 34.3 49.14 73.53 83.99 33.62 33.62 33.62 33.62 28.74 35.32 43.45 46.94
Avg. 17.16 21.05 21.24 21.24 32.07 40.01 49.52 52.13 17.97 19.05 20.08 20.08 22.40 26.70 30.28 31.15
Table 1: Percent (%) improvement in stability loss $\mathcal{L}_{S}$ as function of the accuracy tolerance $\alpha$ across models and datasets.
Dataset Yearbook Healthcare Huffpost Avg.
$\alpha$ 0.01 0.02 0.05 0.1 0.01 0.02 0.05 0.1 0.01 0.02 0.05 0.1 0.01 0.02 0.05 0.1
LogReg 0.66 0.66 0.66 0.66 0.50 0.70 0.89 0.89 0.27 0.27 0.27 0.27 0.48 0.54 0.61 0.61
CART 0.43 2.15 1.57 1.57 -0.02 0.31 0.20 0.20 1.29 0.78 2.98 2.98 0.56 1.08 1.58 1.58
XGBoost -0.05 -0.05 -0.05 -0.05 -0.07 0.03 0.03 0.03 0.21 0.21 0.21 0.21 0.03 0.06 0.06 0.06
MLP 0.31 1.18 1.18 1.18 1.44 -4.51 -0.43 15.01 0.19 0.19 0.19 0.19 0.65 -1.05 0.31 4.13
Avg. 0.34 0.99 0.84 0.84 0.46 -0.87 0.17 4.03 0.49 0.36 0.91 0.91 0.43 0.16 0.64 1.93
Table 2: Percent (%) deterioration of realized out-of-sample accuracy as a function of the selected $\alpha$ across models and datasets.
Robustness check: inter-sequence stability.
@@
Table 3: Inter-sequence stability loss for varying sequence lengths
@@
Next, focusing on the healthcare case study, we empirically investigate the sensitivity of SVML to different update frequencies by splitting the available 52 months of data into intervals of different lengths (1, 3, 6, and 9 months). As discussed, the currently deployed 6-month frequency will change, so our experiment also serves to understand the properties of different frequencies. We choose XGBoost as the canonical model type for the experiment since this is the model in use in production, and train 25 candidate models per batch. We find that trends and insights are consistent across the different frequencies tested. A 3-month re-training frequency exhibits more stable behavior compared to 1-month, while retaining high AUC ranging from 0.85-0.9 with improvements in distance from [0.3-0.37] to [0.275-0.21] compared to 6- or 9-month intervals. We present the detailed results in Appendix D.
@@
As a final robustness check, we investigate how well the analytical insights of an SVML sequence optimized for one distance measure (XGBoost with Gain importance) generalize to a different measure (SHAP) in terms of their stability across retraining iterations. Specifically, we study the empirical distribution of the absolute deviation in features’ SHAP values between adjacent models for both greedy and SVML sequences. Greedy achieves a mean deviation of 0.086 with a standard deviation of 0.101 and a maximum deviation of 0.625; in comparison, SVML exhibits an improved mean deviation of 0.081 with a standard deviation of 0.094 and a maximum deviation of 0.606. These results indicate that SVML’s improved stability across retraining iterations effectively translates into consistent SHAP values, affirming the method’s reliability. We present the detailed results in Appendix D.